A radio telescope with a diameter of 310m uses a wavelength of 5.0m to observe a binary star system that is about 2.5
Answer #1

d = 1.22*5*2.5 x 10^21/(310) = 4.9 x 10^19 m = 4.9 x 10^16 km
Answer #2

The minimum distance of two stars that can be distinguished by the telescope is,

            d = 1.22(lambda)(D) / a

               = (1.22)(5.0 m)(2.5x1021 m) / (310 m)

               = 4.92x1019 m

               = 4.92x1016 km